A non-relativistic proton beam passes without deviation through the region of space where there are uniform transverse mutually perpendicular electric and magnetic fields- with E=120kV/m and B=50mT. Then the beam strikes a grounded target. Find the force with which the beam acts on the target if the beam current I=0.80mA


Grade:upto college level

1 Answers

ROSHAN MUJEEB
3 years ago

Fe​=Fm​
oreE=eBv
∴v=BE​=50×10−3120×103​=2.4×106m/s
Let n be the number of protons striking per second.Then,
ne=0.8×10−3
orn=1.6×10−190.8×10−3​=5×1015/s
Force imparted =rate of change of momentum
=nmv
=5×1015×1.67×10−27×2.4×106
=2.0×10−5N